# David Martínez-Crespo1,2, Cesare Tronci3 1Departamento de Física Teórica, Universidad de Zaragoza, Zaragoza, Spain 2Centro de Astropartículas y Física de Altas Energías, Universidad de Zaragoza, Zaragoza, Spain

### v1

@@ 3 Heisenberg picture of hybrid Koopman dynamics
$\Big{[}\hat{\rho}_{0},i\hbar(\hat{\zeta}+\boldsymbol{\cal X}\cdot\widehat{\boldsymbol{\gamma}})-{\widehat{\mathcal{H}}}-i\hbar\big{[}{\rm d}^{\widehat{{\gamma}}}\hat{\rho}_{0}+\hat{\rho}_{0}{\nabla\ln\sqrt{f_{0}}},\,\widehat{\!\boldsymbol{\cal Z}\,}\!\big{]}\Big{]}=0,$
where the notation is such that $[{\rm d}^{\widehat{{\gamma}}}\hat{\rho}_{0}+\hat{\rho}_{0}{\nabla\ln\sqrt{f_{0}}},\,\widehat{\!\boldsymbol{\cal Z}\,}\!]=[\partial_{k}^{\widehat{{\gamma}}}\hat{\rho}_{0}+\hat{\rho}_{0}{\partial_{k}\ln\sqrt{f_{0}}},\widehat{\cal Z}^{k}]$. Upon recalling $\,\widehat{\!\boldsymbol{\cal Z}\,}\!={{\boldsymbol{\mathcal{X}}}}_{\widehat{\cal H}}$, this implies
@@ 3.1 The convective vector field
$-\frac{1}{2}\langle\nabla^{\widehat{{\gamma}}}\hat{\rho}_{0},i\hbar[\hat{\rho}_{0},\pounds_{\,\widehat{\!\boldsymbol{\cal Z}\,}\!}\ln f_{0}]\rangle$
$=$ $\,\langle\hat{\rho}_{0},\nabla^{\widehat{{\gamma}}}{\widehat{\mathcal{H}}}\rangle+\frac{\hbar}{2}\big{(}\langle\pounds_{\,\widehat{\!\boldsymbol{\cal Z}\,}\!},\widehat{\boldsymbol{\Gamma}}\rangle+\langle\widehat{\cal Z}^{k},[{\widehat{\gamma}}_{k},\widehat{\boldsymbol{\Gamma}}]\rangle-\langle[\widehat{\cal Z}^{k},\widehat{\Gamma}_{k}],{\widehat{\boldsymbol{\gamma}}}\rangle+\langle\widehat{\boldsymbol{\Gamma}},\pounds_{\,\widehat{\!\boldsymbol{\cal Z}\,}\!}\ln f_{0}\rangle\big{)}$
where $\widehat{\boldsymbol{\Gamma}}=i[\hat{\rho}_{0},\nabla^{\widehat{{\gamma}}}\hat{\rho}_{0}]$ and we used the covariant Leibniz rule $\nabla\langle\widehat{A},\widehat{B}\rangle=\langle\nabla^{\widehat{{\gamma}}}\widehat{A},\widehat{B}\rangle+\langle\widehat{A},\nabla^{\widehat{{\gamma}}}\widehat{B}\rangle$ as well as the relation $[\partial_{k}^{\widehat{{\gamma}}},\partial_{h}^{\widehat{{\gamma}}}]=0$, which follows from the fact that $\widehat{\boldsymbol{\gamma}}$ is a flat connection. As before, the Lie derivative notation is such that, for example, £$\langle\pounds_{\widehat{\cal Z}},\widehat{\Gamma}\rangle=\langle\pounds_{\,\widehat{\!\boldsymbol{\cal Z}\,}\!},\widehat{\boldsymbol{\Gamma}}\rangle\cdot{\rm d}{\mathbf{z}}$ and £$\pounds_{\widehat{\cal Z}}\ln f_{0}=\pounds_{\,\widehat{\!\boldsymbol{\cal Z}\,}\!}\ln f_{0}$. Consequently, upon introducing the Poisson tensor (23), we obtain
$\boldsymbol{\cal X}=\langle\hat{\rho}_{0},\,\widehat{\!\boldsymbol{\cal Z}\,}\!\rangle+\frac{\hbar}{2}\Big{(}\langle\Pi\widehat{\boldsymbol{\Gamma}},\pounds_{\,\widehat{\!\boldsymbol{\cal Z}\,}\!}\ln f_{0}\rangle+\langle\pounds_{\,\widehat{\!\boldsymbol{\cal Z}\,}\!},\Pi\widehat{\boldsymbol{\Gamma}}\rangle-\langle(\pounds_{\,\widehat{\!\boldsymbol{\cal Z}\,}\!}\Pi+[\widehat{{\gamma}}_{k},\widehat{\cal Z}^{k}]\Pi),\widehat{\boldsymbol{\Gamma}}\rangle-\langle[\widehat{\cal Z}^{k},\widehat{\Gamma}_{k}],\Pi{\widehat{\boldsymbol{\gamma}}}\rangle\Big{)}$
@@
where $(\boldsymbol{\cal X}_{{\widehat{\mathcal{H}}}}\wedge\widehat{\boldsymbol{\Xi}})^{jk}={\cal X}_{{\widehat{\mathcal{H}}}}^{j}\widehat{\Xi}^{k}-\widehat{\Xi}^{j}{\cal X}_{{\widehat{\mathcal{H}}}}^{k}$ is a bivector. Also, here the second equality follows from (38) upon using £$\pounds_{\boldsymbol{\cal X}_{{\widehat{\mathcal{H}}}}}^{\widehat{\gamma}}\Pi=0$ and
@@
^$\frac{\partial{\widehat{\mathcal{H}}}}{\partial t}=$ $\ \Pi\Big{(}{\rm d}^{\widehat{{\gamma}}}\widehat{\mathcal{H}},{\big{\langle}{\rm d}^{\widehat{{\gamma}}}\widehat{\mathcal{H}}\big{\rangle}}+\frac{\hbar}{2f_{0}}\big{\langle}\pounds_{{\cal X}_{{\widehat{\mathcal{H}}}}}^{\widehat{\gamma}},f_{0}\widehat{\Gamma}\big{\rangle}\Big{)}-\frac{1}{2}\Big{[}\big{[}2{\rm d}^{\widehat{{\gamma}}}\hat{\rho}_{0}+\hat{\rho}_{0}{\rm d}\ln{f_{0}},{\cal X}_{\widehat{\cal H}}\big{]},{\widehat{\mathcal{H}}}\Big{]},$ (40)
@@
where, in coordinates, £$\langle\pounds_{\boldsymbol{\cal X}_{{\widehat{\mathcal{H}}}}}^{\widehat{\gamma}},f_{0}\widehat{\boldsymbol{\Gamma}}\rangle=\langle\partial_{k}^{\widehat{\gamma}}{{\cal X}_{{\widehat{\mathcal{H}}}}}^{k},f_{0}\widehat{\boldsymbol{\Gamma}}\rangle+\langle{{\cal X}_{{\widehat{\mathcal{H}}}}}^{k},\partial_{k}^{\widehat{\gamma}}f_{0}\widehat{\boldsymbol{\Gamma}}\rangle+\langle f_{0}\widehat{\Gamma}_{k},\nabla^{\widehat{\gamma}}{{\cal X}_{{\widehat{\mathcal{H}}}}}^{k}\rangle$ and ${\rm d}^{\widehat{\gamma}}\widehat{\Gamma}=i[{\rm d}^{\widehat{\gamma}}\hat{\rho}_{0},\wedge{\rm d}^{\widehat{\gamma}}\hat{\rho}_{0}]$. Once again, we also recall the definition $\Pi=-\Omega^{-1}$ as in (23).
The formidable appearance of these equations reflects the intricate nature of the quantum backreaction, which is responsible for the levels of complexity already appeared in the Schrödinger picture; see equations (2)-(4). Nevertheless, it is possible to see that, similarly to the case of Ehrenfest dynamics, the purely quantum motion is obtained if ${\rm d}^{\widehat{{\gamma}}}{\widehat{\mathcal{H}}}=0$, so that ${\cal X}_{{\widehat{\mathcal{H}}}}=0$ and the entire system returns trivial dynamics. Similarly, the purely classical motion is again obtained in the case ^${\widehat{\mathcal{H}}}={\cal H}\boldsymbol{1}$, so that equation (41) becomes $i\hbar{\partial_{t}\widehat{{\gamma}}}={\rm d}{\cal H}$, which enforces $\widehat{{\gamma}}=i{\rm d}\phi\boldsymbol{1}$ for some function $\phi$. In this case, ${\rm d}^{\widehat{{\gamma}}}={\rm d}$ and the first two equations again return trivial dynamics.
@@ 3.2 Heisenberg dynamics and expectation values
so that the local expectation value obeys
$\frac{\partial\langle\widehat{\cal O}\rangle}{\partial t}=$ $\ \left\langle\boldsymbol{\cal X}_{{\widehat{\mathcal{H}}}}\cdot\langle\nabla^{\widehat{{\gamma}}}\widehat{{\cal O}}\rangle,\hat{\rho}_{0}+\frac{\hbar}{2f_{0}}\operatorname{div}^{\widehat{\gamma}}\widehat{\boldsymbol{\Xi}}\right\rangle+\frac{\hbar}{2f_{0}}\operatorname{div}\!\big{(}\langle\boldsymbol{\cal X}_{{\widehat{\mathcal{H}}}},\widehat{\boldsymbol{\Xi}}\rangle-\langle\widehat{\boldsymbol{\Xi}},\boldsymbol{\cal X}_{{\widehat{\mathcal{H}}}}\rangle\big{)}\cdot\langle\nabla^{\widehat{{\gamma}}}\widehat{{\cal O}}\rangle$
@@
At this point, it may be revealing to consider the dynamics of the overall expectation value $\int\!f_{0}\langle\widehat{\cal O}\rangle\,{\rm d}^{2}z_{0}$ for a purely classical and quantum observable, respectively ${\cal O}_{C}$ and $\widehat{\cal O}_{Q}$. Upon introducing the notation $\{\widehat{A},\widehat{B}\}^{\widehat{{\gamma}}}=\Pi({\rm d}^{\widehat{{\gamma}}}\widehat{A},{\rm d}^{\widehat{{\gamma}}}\widehat{B})$, we have
$\frac{\partial}{\partial t}\int\!f_{0}{\cal O}_{C}\,{\rm d}^{2}z_{0}=$$\int\!{\Big{\langle}\{{\cal O}_{C},\widehat{\mathcal{H}}\}^{\widehat{{\gamma}}},f_{0}\hat{\rho}_{0}+\frac{\hbar}{2}\operatorname{div}^{\widehat{{\gamma}}}\widehat{\boldsymbol{\Xi}}\Big{\rangle}}\,{\rm d}^{2}z_{0},$
$\frac{\partial}{\partial t}\int\!f_{0}\langle\widehat{\cal O}_{Q}\rangle\,{\rm d}^{2}z_{0}=$ $\int\!\Big{\langle}i\hbar^{-1}[{\widehat{\mathcal{H}}},\widehat{{\cal O}}_{Q}],f_{0}\hat{\rho}_{0}+\frac{\hbar}{2}\operatorname{div}^{\widehat{{\gamma}}}\widehat{\boldsymbol{\Xi}}\Big{\rangle}\,{\rm d}^{2}z_{0},$
where we have used $\langle{\cal O}_{C}\rangle={\cal O}_{C}$ and ${\rm d}^{\widehat{{\gamma}}}\widehat{\cal O}_{Q}=0$, so that $\langle\widehat{\cal O}_{Q},\operatorname{div}^{\widehat{{\gamma}}}\widehat{\boldsymbol{\cal W}}\rangle=\operatorname{div}\langle\widehat{\cal O}_{Q},\widehat{\boldsymbol{\cal W}}\rangle$. Also, the first equation above follows by integration by parts from the skew-symmetry of the bivector-density $\langle\boldsymbol{\cal X}_{{\widehat{\mathcal{H}}}},\widehat{\boldsymbol{\Xi}}\rangle-\langle\widehat{\boldsymbol{\Xi}},\boldsymbol{\cal X}_{{\widehat{\mathcal{H}}}}\rangle$. Notice that classical and quantum conservation laws follow immediately in the case $\{{\cal O}_{C},\widehat{\mathcal{H}}\}^{\widehat{{\gamma}}}=0$ and [^$[{\widehat{\mathcal{H}}},\widehat{{\cal O}}_{Q}]=0$, respectively. Importantly, we observe the distinctive contribution of the backreaction forces, encoded by the divergence terms, in both evolutions laws. These terms indicate substantial differences between the Ehrenfest and Koopman models.
While the expectation value dynamics above offers some insight, the intricate structure of the terms associated to the quantum backreaction leaves little room to physical intuition. At present, the role of these terms in mixed quantum-classical Heisenberg dynamics can hardly be understood without resorting to suitable examples. In particular, here we want to compare the Ehrenfest and the Koopman models for a specific problem.

### v2

@@ 3 Heisenberg picture of hybrid Koopman dynamics
$\Big{[}\hat{\rho}_{0},i\hbar(\hat{\zeta}+\boldsymbol{\cal X}\cdot\widehat{\boldsymbol{\gamma}})-{\widehat{\mathcal{H}}}-i\hbar\big{[}\nabla^{\widehat{{\gamma}}}\hat{\rho}_{0}+\hat{\rho}_{0}{\nabla{\ln}\sqrt{f_{0}}},\,\widehat{\!\boldsymbol{\cal Z}\,}\!\big{]}\Big{]}=0,$
where the notation is such that $[\nabla^{\widehat{{\gamma}}}\hat{\rho}_{0}+\hat{\rho}_{0}{\nabla{\ln}\sqrt{f_{0}}},\,\widehat{\!\boldsymbol{\cal Z}\,}\!]=[\partial_{k}^{\widehat{{\gamma}}}\hat{\rho}_{0}+\hat{\rho}_{0}{\partial_{k}{\ln}\sqrt{f_{0}}},\widehat{\cal Z}^{k}]$. Upon recalling $\,\widehat{\!\boldsymbol{\cal Z}\,}\!={{\boldsymbol{\mathcal{X}}}}_{\widehat{\cal H}}$, this implies
@@ 3.1 The convective vector field
$-\frac{1}{2}\langle\nabla^{\widehat{{\gamma}}}\hat{\rho}_{0},i\hbar[\hat{\rho}_{0},\pounds_{\widehat{\cal Z}}\ln f_{0}]\rangle$
$=$ $\,\langle\hat{\rho}_{0},\nabla^{\widehat{{\gamma}}}{\widehat{\mathcal{H}}}\rangle+\frac{\hbar}{2}\big{(}\langle\pounds_{\,\widehat{\!\boldsymbol{\cal Z}\,}\!},\widehat{\boldsymbol{\Gamma}}\rangle+\langle\widehat{\cal Z}^{k},[{\widehat{\gamma}}_{k},\widehat{\boldsymbol{\Gamma}}]\rangle-\langle[\widehat{\cal Z}^{k},\widehat{\Gamma}_{k}],{\widehat{\boldsymbol{\gamma}}}\rangle+\langle\widehat{\boldsymbol{\Gamma}},\pounds_{\widehat{\cal Z}}\ln f_{0}\rangle\big{)}$
where $\widehat{\boldsymbol{\Gamma}}=i[\hat{\rho}_{0},\nabla^{\widehat{{\gamma}}}\hat{\rho}_{0}]$ and we used the covariant Leibniz rule $\nabla\langle\widehat{A},\widehat{B}\rangle=\langle\nabla^{\widehat{{\gamma}}}\widehat{A},\widehat{B}\rangle+\langle\widehat{A},\nabla^{\widehat{{\gamma}}}\widehat{B}\rangle$ as well as the relation $[\partial_{k}^{\widehat{{\gamma}}},\partial_{h}^{\widehat{{\gamma}}}]=0$, which follows from the fact that $\widehat{\boldsymbol{\gamma}}$ is a flat connection. As before, the Lie derivative notation is such that, for example, £$\langle\pounds_{\widehat{\cal Z}},\widehat{\Gamma}\rangle=\langle\pounds_{\,\widehat{\!\boldsymbol{\cal Z}\,}\!},\widehat{\boldsymbol{\Gamma}}\rangle\cdot{\rm d}{\mathbf{z}}$ and £$\pounds_{\widehat{\cal Z}}\ln f_{0}=\,\widehat{\!\boldsymbol{\cal Z}\,}\!\,\lrcorner\,{{\rm d}\ln}f_{0}=\,\widehat{\!\boldsymbol{\cal Z}\,}\!\cdot\nabla{\ln}f_{0}$. Consequently, upon introducing the Poisson tensor (23), we obtain
$\boldsymbol{\cal X}=\langle\hat{\rho}_{0},\,\widehat{\!\boldsymbol{\cal Z}\,}\!\rangle+\frac{\hbar}{2}\Big{(}\langle\Pi\widehat{\boldsymbol{\Gamma}},\pounds_{\widehat{\cal Z}}\ln f_{0}\rangle+\langle\pounds_{\,\widehat{\!\boldsymbol{\cal Z}\,}\!},\Pi\widehat{\boldsymbol{\Gamma}}\rangle-\langle(\pounds_{\,\widehat{\!\boldsymbol{\cal Z}\,}\!}\Pi+[\widehat{{\gamma}}_{k},\widehat{\cal Z}^{k}]\Pi),\widehat{\boldsymbol{\Gamma}}\rangle-\langle[\widehat{\cal Z}^{k},\widehat{\Gamma}_{k}],\Pi{\widehat{\boldsymbol{\gamma}}}\rangle\Big{)}$
@@
where $(\boldsymbol{\cal X}_{{\widehat{\mathcal{H}}}}\wedge\widehat{\boldsymbol{\Xi}})^{jk}={\cal X}_{{\widehat{\mathcal{H}}}}^{j}\widehat{\Xi}^{k}-\widehat{\Xi}^{j}{\cal X}_{{\widehat{\mathcal{H}}}}^{k}$ is a bivector density. Also, here the second equality follows from (38) upon using £$\pounds_{\boldsymbol{\cal X}_{{\widehat{\mathcal{H}}}}}^{\widehat{\gamma}}\Pi=0$ and
@@
^$\frac{\partial{\widehat{\mathcal{H}}}}{\partial t}=$ $\ \Pi\Big{(}{\rm d}^{\widehat{{\gamma}}}\widehat{\mathcal{H}},{\big{\langle}{\rm d}^{\widehat{{\gamma}}}\widehat{\mathcal{H}}\big{\rangle}}+\frac{\hbar}{2f_{0}}\big{\langle}\pounds_{{\cal X}_{{\widehat{\mathcal{H}}}}}^{\widehat{\gamma}},f_{0}\widehat{\Gamma}\big{\rangle}\Big{)}-\Big{[}\big{[}{{\rm d}^{\widehat{{\gamma}}}\hat{\rho}_{0}+\hat{\rho}_{0}{{\rm d}\ln}\sqrt{{f_{0}}}},{\cal X}_{\widehat{\cal H}}\big{]},{\widehat{\mathcal{H}}}\Big{]},$ (40)
@@
where ${\rm d}^{\widehat{\gamma}}\widehat{\Gamma}=i[{\rm d}^{\widehat{\gamma}}\hat{\rho}_{0},\wedge{\rm d}^{\widehat{\gamma}}\hat{\rho}_{0}]$ and, in coordinates, £$\langle\pounds_{\boldsymbol{\cal X}_{{\widehat{\mathcal{H}}}}}^{\widehat{\gamma}},f_{0}\widehat{\boldsymbol{\Gamma}}\rangle=\langle\partial_{k}^{\widehat{\gamma}}{{\cal X}_{{\widehat{\mathcal{H}}}}}^{k},f_{0}\widehat{\boldsymbol{\Gamma}}\rangle+\langle{{\cal X}_{{\widehat{\mathcal{H}}}}}^{k},\partial_{k}^{\widehat{\gamma}}f_{0}\widehat{\boldsymbol{\Gamma}}\rangle+\langle f_{0}\widehat{\Gamma}_{k},\nabla^{\widehat{\gamma}}{{\cal X}_{{\widehat{\mathcal{H}}}}}^{k}\rangle$. Once again, we also recall the definition $\Pi=-\Omega^{-1}$ as in (23).
The formidable appearance of these equations reflects the intricate nature of the quantum backreaction, which is responsible for the levels of complexity already appeared in the Schrödinger picture; see equations (2)-(4). Nevertheless, it is possible to see that, similarly to the case of Ehrenfest dynamics, the purely quantum motion is obtained if ${\rm d}^{\widehat{{\gamma}}}{\widehat{\mathcal{H}}}=0$, so that ${\cal X}_{{\widehat{\mathcal{H}}}}=0$ and the entire system returns trivial dynamics. Similarly, the purely classical motion is again obtained in the case ^${\widehat{\mathcal{H}}}={\cal H}\boldsymbol{1}$, so that equation (41) becomes $i\hbar{\partial_{t}\widehat{{\gamma}}}={\rm d}{\cal H}$, which enforces $\widehat{{\gamma}}=i{\rm d}\phi\boldsymbol{1}$ for some function $\phi$ such that ${\rm d}\phi|_{t=0}=0$. In this case, ${\rm d}^{\widehat{{\gamma}}}={\rm d}$ and the first two equations again return trivial dynamics.
We also emphasize that, although in this case the local expectation $\langle\widehat{\cal H}\rangle$ is not conserved, conservation of the total energy in (8) is guaranteed by the Hamiltonian structure of the Koopman model. In the Heisenberg picture, this total energy is expressed as $h(\widehat{\mathcal{H}},\widehat{\gamma})={\int}f_{0}\langle\widehat{\cal H}+i\hbar[\nabla^{\widehat{\gamma}}\hat{\rho}_{0},{\boldsymbol{\mathcal{X}}}_{{\widehat{\mathcal{H}}}}]/2\rangle{\rm d}^{2}z_{0}$. As mentioned previously, here the second term is responsible for generating correlations between the quantum and the classical system in interaction.
@@ 3.2 Heisenberg dynamics and expectation values
so that, switching back to coordinate notation, the local expectation value obeys
$\frac{\partial\langle\widehat{\cal O}\rangle}{\partial t}=$ $\ \left\langle\boldsymbol{\cal X}_{{\widehat{\mathcal{H}}}}\cdot\langle\nabla^{\widehat{{\gamma}}}\widehat{{\cal O}}\rangle,\hat{\rho}_{0}+\frac{\hbar}{2f_{0}}\operatorname{div}^{\widehat{\gamma}}\widehat{\boldsymbol{\Xi}}\right\rangle+\frac{\hbar}{2f_{0}}\operatorname{div}{\operatorname{Tr}\!\big{(}\boldsymbol{\cal X}_{{\widehat{\mathcal{H}}}}\wedge\widehat{\boldsymbol{\Xi}}\big{)}}\cdot\langle\nabla^{\widehat{{\gamma}}}\widehat{{\cal O}}\rangle$
@@
At this point, it may be revealing to consider the dynamics of the overall expectation value $\int\!f_{0}\langle\widehat{\cal O}\rangle\,{\rm d}^{2}z_{0}$ for a purely classical and quantum observable, respectively ${\cal O}_{\sf C}$ and $\widehat{\cal O}_{\sf Q}$. Upon introducing the notation $\{\widehat{A},\widehat{B}\}^{\widehat{{\gamma}}}=\Pi({\rm d}^{\widehat{{\gamma}}}\widehat{A},{\rm d}^{\widehat{{\gamma}}}\widehat{B})$, we have
$\frac{\partial}{\partial t}\int\!f_{0}{\cal O}_{\sf C}\,{\rm d}^{2}z_{0}=$$\int\!{\Big{\langle}\{{\cal O}_{\sf C},\widehat{\mathcal{H}}\}^{\widehat{{\gamma}}},f_{0}\hat{\rho}_{0}+\frac{\hbar}{2}\operatorname{div}^{\widehat{{\gamma}}}\widehat{\boldsymbol{\Xi}}\Big{\rangle}}\,{\rm d}^{2}z_{0},$
$\frac{\partial}{\partial t}\int\!f_{0}\langle\widehat{\cal O}_{\sf Q}\rangle\,{\rm d}^{2}z_{0}=$ $\int\!\Big{\langle}i\hbar^{-1}[{\widehat{\mathcal{H}}},\widehat{{\cal O}}_{\sf Q}],f_{0}\hat{\rho}_{0}+\frac{\hbar}{2}\operatorname{div}^{\widehat{{\gamma}}}\widehat{\boldsymbol{\Xi}}\Big{\rangle}\,{\rm d}^{2}z_{0},$
where we have used $\langle{\cal O}_{\sf C}\rangle={\cal O}_{\sf C}$ and ${\rm d}^{\widehat{{\gamma}}}\widehat{\cal O}_{\sf Q}=0$, so that $\langle\widehat{\cal O}_{\sf Q},\operatorname{div}^{\widehat{{\gamma}}}\widehat{\boldsymbol{\cal W}}\rangle=\operatorname{div}\langle\widehat{\cal O}_{\sf Q},\widehat{\boldsymbol{\cal W}}\rangle$. Also, the first equation above follows by integration by parts from the skew-symmetry of the bivector-density $\operatorname{Tr}\!\big{(}\boldsymbol{\cal X}_{{\widehat{\mathcal{H}}}}\wedge\widehat{\boldsymbol{\Xi}}\big{)}$. Notice that classical and quantum conservation laws follow immediately in the case $\{{\cal O}_{\sf C},\widehat{\mathcal{H}}\}^{\widehat{{\gamma}}}=0$ and [^$[{\widehat{\mathcal{H}}},\widehat{{\cal O}}_{\sf Q}]=0$, respectively. Importantly, we observe the distinctive contribution of the backreaction forces, encoded by the divergence terms, in both evolutions laws. These terms indicate substantial differences between the Ehrenfest and Koopman models. For example, consider the translation invariant Hamiltonian ^$\widehat{\mathcal{H}}=(M^{-1}{\mathcal{P}}^{2}+m^{-1}\widehat{\!\mathscr{K}\,}\!^{2})/2+V(\mathcal{Q}-\widehat{\!\mathscr{X}})$ at the end of Remark 5: as discussed, the Ehrenfest dynamics conserves the local expectation ^$\langle\widehat{\mathcal{M}}\rangle={\mathcal{P}}+\langle\,\widehat{\!\mathscr{K}}\rangle$, while Koopman dynamics only conserves the global momentum $\int\!f_{0}({\mathcal{P}}+\langle\,\widehat{\!\mathscr{K}}\rangle)\,{\rm d}^{2}z_{0}$. Indeed, since ${\rm d}^{\widehat{{\gamma}}\,}\widehat{\!\mathscr{X}}=0$, we have ${\rm d}^{\widehat{{\gamma}}}V=\partial_{\mathcal{Q}}V\,{\rm d}\mathcal{Q}$ and $\{\mathcal{P},\widehat{\mathcal{H}}\}^{\widehat{{\gamma}}}=i\hbar^{-1}[\,\widehat{\!\mathscr{K}},{\widehat{\mathcal{H}}}]$. We notice that the total momentum conservation may be related to Galilean invariance in the hybrid quantum-classical setting [9], although here we leave this aspect for future work.
While the expectation value dynamics offers some insight, the intricate structure of the terms associated to the quantum backreaction leaves little room to physical intuition. At present, the role of these terms in mixed quantum-classical Heisenberg dynamics can hardly be understood without resorting to suitable examples. In particular, here we want to compare the Ehrenfest and the Koopman models for a specific problem.